# Stock Price Responses to Firm-Level News in Supply Chain Networks

### v4

@@ 5 Discussion
In summary, we find that the way positive and negative information about firms affects stock prices of their suppliers and clients before and after the disclosure of the information by news media is different between firms in Japan and those in other countries. Additionally, we find that stronger pre-news stock price responses are observed for positive news in the Japanese sample.
It is important to emphasize that these results do not imply that investors possess detailed knowledge of firm-level supply chain structures. Rather, the observed stock price responses are consistent with market pricing that reflects information correlated with supply chain relationships. Several factors may be correlated with the observed stock price responses along supply chain relationships. The present analysis does not identify the mechanisms underlying these correlations.
It is also important to distinguish between realized economic impacts and market-based expectations reflected in stock prices. The present study focuses on stock price responses to news, which capture investors’ expectations about potential spillovers along supply chain relationships rather than confirmed changes in production, orders, or real economic outcomes. Whether such expectations are eventually realized in the form of actual production disruptions or demand reductions is beyond the scope of this analysis.
@@
In this context, we clarify the specific dimension of market responses examined in this study. The focus of this study is on stock price responses, which reflect changes in market valuation. While other market indicators such as trading volume or volatility capture important aspects of market activity and uncertainty, they represent different dimensions of market behavior and are not direct measures of valuation effects. Accordingly, the present analysis concentrates on price-based responses to quantify how news-related information propagates through supply chain networks. Analyses of news spillovers in trading volume or volatility would require a different empirical design and are therefore not pursued in this study.
With this focus on price-based responses in mind, we turn to the interpretation of differences between the global and Japanese samples. These differences between the global and Japanese samples do not imply that different databases mechanically generate different results. Rather, the results consistently show that stock price responses propagate through supply chain relationships in both samples, while the timing and relative magnitude of such propagation differ across institutional contexts.
The use of different supply chain databases reflects these institutional contexts rather than causing the observed differences. The FactSet database captures major supply chain links disclosed in financial statements, whereas the TSR database identifies economically important trading partners through firm-level surveys. These databases therefore provide complementary views of supply chain relationships, rather than alternative measurements of the same underlying network.
Although the FactSet database is mainly determined by criteria based on sales volume, the TSR database identifies important trade partners, which are not determined only by sales volume. One possible interpretation is that the observed differences are consistent with heterogeneous timing and magnitude of market responses across institutional contexts. The present analysis does not establish the mechanisms generating these patterns.
Several institutional and historical factors may potentially contribute to the observed differences between Japan and other countries. The observed differences arise in an institutional context in which supply chain relationships in Japan are often discussed using the term keiretsu [3, 50]. The present analysis does not identify whether, or how, such institutional features contribute to the observed patterns. Other factors, such as historical shocks to domestic supply chains or differences between domestic and global supply chain structures, may also play a role. Identifying the relative importance of these mechanisms is beyond the scope of this study and remains an important direction for future research.
@@ 6 Conclusion
We find that news sentiment, both positive and negative, is associated with systematically higher and lower change rates of stock prices of firms mentioned in the news in the pre-news period, which may be consistent with a combination of anticipatory market behavior, information correlated with forthcoming news, and selection in news coverage. Such news sentiment is also associated with stock price responses of the firms’ suppliers and clients in both the pre-news and post-news periods, consistent with systematic valuation comovements along supply-chain relationships. In addition, we generally find that post-news effects on stock prices of the mentioned firms and their suppliers and clients are larger than pre-news effects. The difference between post- and pre-news effects summarizes differences in stock price changes between the post- and pre-disclosure periods. However, for Japanese firms, post-news effects on suppliers and clients can be smaller than pre-news effects, in contrast to the pattern observed for firms across the world.

### v5

@@ 5 Discussion
In summary, we find that positive and negative news about firms are associated with stock price changes of their suppliers and clients before and after the disclosure of the information and the pattern of association differs between the world and Japan, which is observed by large scale data of news, stock prices, and supply chains. Additionally, we find that stronger pre-news stock price responses are observed for positive news in the Japanese sample.
As noted in the Introduction, we adopt a reduced-form perspective in the sense that we do not model investors’ information processing or decision-making. With this focus on price-based responses in mind, we turn to the interpretation of differences between the global and Japanese samples. The differences between the global and Japanese samples do not imply that different databases mechanically generate different results. Rather, the results consistently show that stock price responses suppliers’ and clients’ news in both samples, while the timing and relative magnitude of association differ across global and Japanese contexts. The FactSet database captures major supply chain links disclosed in financial statements, whereas the TSR database identifies economically important trading partners through firm-level surveys. These databases therefore provide qualitatively different views of supply chain relationships. The specific context of Japan may potentially contribute to the observed differences. Supply chain relationships in Japan are often discussed using the term keiretsu [4, 50]. Such institutional difference may possibly affect the decision making of investors. The present analysis does not identify whether, or how, such institutional features contribute to the observed patterns. Other factors, such as historical shocks to Japanese supply chains due to disasters, such as the Great East Japan earthquake in 2011, may also play a role. Again, identifying the difference behind supply chains is beyond the scope of this study and remains an important direction for future research.
@@
We examine the stock price responses, i.e., market valuation. While other market indicators such as trading volume or volatility capture important aspects of market activity and uncertainty, they represent different dimensions of market behavior. Analyses of news spillovers in trading volume or volatility would require a different empirical design and are therefore not pursued in this study.
@@ 6 Conclusion
We find that news sentiment, both positive and negative, is associated with systematically higher and lower change rates of stock prices of firms mentioned in the news in the pre- and post-news period. News sentiment is also associated with stock price responses of the firms’ suppliers and clients in both the pre-news and post-news periods, consistent with systematic valuation comovements along supply-chain relationships observed in other studies. In addition, we generally find that post-news effects on stock prices of the mentioned firms and their suppliers and clients are larger than pre-news effects. However, for Japanese firms, post-news effects on suppliers and clients can be smaller than pre-news effects, in contrast to the pattern observed for firms across the world.